OpenStudy (nannie):

Please show me how to Rationalize the Denominators, assuming all the numbers under ther radicals are positive numbers. √x-√y/√x+√y

OpenStudy (saifoo.khan):

\[\frac{x-2\sqrt{xy}+y}{x-y}\]

OpenStudy (nannie):

What's with you just giving answers?

OpenStudy (anonymous):

\[\frac{\sqrt{x}-\sqrt{y}}{\sqrt{x}+\sqrt{y}}\] \[\frac{\sqrt{x}-\sqrt{y}}{\sqrt{x}+\sqrt{y}}\times \frac{\sqrt{x}-\sqrt{y}}{\sqrt{x}-\sqrt{y}}\] \[\frac{x-y-2\sqrt{xy}}{x-y}\]

OpenStudy (saifoo.khan):

is that -y in the numerator @sat?

OpenStudy (saifoo.khan):

LOL @ nanie. that's a secret.

OpenStudy (anonymous):

nah it is wrong i made a mistake

OpenStudy (nannie):

Yes

OpenStudy (nannie):

And thanks, this is finals week!

OpenStudy (anonymous):

\[\frac{x+y-2\sqrt{xy}}{x-y}\]

OpenStudy (saifoo.khan):

that's right now! :D

OpenStudy (anonymous):

why on earth are finals in october?

OpenStudy (nannie):

Want to finish the rest? I only have 16 left. Finals for the term

OpenStudy (anonymous):

lol post away, might as well get them over with

OpenStudy (saifoo.khan):

LOL.

OpenStudy (nannie):

express using a positive exponents. y^-5

OpenStudy (saifoo.khan):

\[\frac{1}{y^5}\]

OpenStudy (anonymous):

yay

OpenStudy (saifoo.khan):

xD

OpenStudy (nannie):

This I am checking the answer on. Factor completely (7x+2)(9x^2+9x+2)

OpenStudy (saifoo.khan):

isnt it factored?

OpenStudy (saifoo.khan):

sat: explain why it is necessary to have common denominators when adding or subtracting rational expressions.

OpenStudy (nannie):

What kind of calculator do you have.

OpenStudy (saifoo.khan):

Mind calculator! :D

OpenStudy (nannie):

says multiply and simplify

OpenStudy (anonymous):

mental one, also and abacus, but i rarely use is

OpenStudy (saifoo.khan):

i will solve this on Abacus.

OpenStudy (nannie):

Right!!

OpenStudy (saifoo.khan):

\[63 x^3+81 x^2+32 x+4 \]

OpenStudy (anonymous):

\[(3 x+1) (3 x+2) (7 x+2)\]

OpenStudy (anonymous):

said "factor" right?

OpenStudy (nannie):

Yep

OpenStudy (saifoo.khan):

no it said: says multiply and simplify

OpenStudy (anonymous):

This I am checking the answer on. Factor completely (7x+2)(9x^2+9x+2)

OpenStudy (nannie):

thats right Sorry

OpenStudy (saifoo.khan):

Sat, please answer mine as well. :(

OpenStudy (anonymous):

multiply, or factor?

OpenStudy (anonymous):

@saifoo what question? why do you need a common denominator? for the same reason you need a common denominator when you add \[\frac{1}{2}+\frac{1}{3}\]

OpenStudy (nannie):

multiply and simplify

OpenStudy (saifoo.khan):

lol, that is not the right answer man.

OpenStudy (nannie):

Factor this 16c²-49

OpenStudy (saifoo.khan):

\[(4 c-7) (4 c+7) \]

OpenStudy (saifoo.khan):

lol, we are doing all your mathlab hw.

OpenStudy (nannie):

I got -1.75, 1.75?

OpenStudy (saifoo.khan):

wrong, type exactly how i wrote, you will great it right.